OpenStudy (anonymous):

Simplify 3/(x+2) - 5/(x-2)

OpenStudy (anonymous):

no, just simplifying the expression. no equation.

OpenStudy (anonymous):

x = {-1, 3}

OpenStudy (anonymous):

\[\frac{ 3 }{ x+2 } - \frac{ 5 }{ x-2 }\]

OpenStudy (anonymous):

Not solving, simplifying.

OpenStudy (anonymous):

subtracting?

OpenStudy (anonymous):

trying to SIMPLIFY the expression, which i think means get the x out of the denominators, but no equations.

OpenStudy (triciaal):

it basically means find the common denominator and make only one fraction

OpenStudy (anonymous):

how?

OpenStudy (triciaal):

for example if you have 1/3 of a pizza and 2/5 of a pizza how much pizza do you really have?

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bc}\] is the one true way to subtract fractions do that here

OpenStudy (anonymous):

you will not get the x out of the denominator they stay there

OpenStudy (triciaal):